OpenStudy (anonymous):

A 10.0g sample of XCl2 reacts with a sufficient amount of Cl2 to convert all of the XCl2 into 12.55g of XCl4. What is the atomic weight of X?

OpenStudy (xishem):

\[X+2(Cl)=10.0g\]\[X+4(Cl)=12.55g\]Now let's just use elimination to combine these two equations...\[-2X-4(Cl)=-20.0g\]\[X+4(Cl)=12.55g\]\[-X=-7.5g \rightarrow X=7.5g\]Therefore, the atomic weight of X is 7.5g.

OpenStudy (anonymous):

Thanks, that makes sense